OpenStudy (anonymous):

what it the missing interger to: ? - (-2) = -6

OpenStudy (mathstudent55):

x - (-2) = -6 x + 2 = -6 x = -8

OpenStudy (anonymous):

?-(-2)=-6 ?+2=-6 ?+2-2=-6-2 ?=-6-2